Dimensional analysis is the most powerful tool for solving word problems. Since chemistry is all word problems, you will see many dimensional analysis practice problems to help strengthen your problem-solving skills. Use the made-up nursery rhyme example to practice dimensional analysis.

"Humpty Dumpty sat on a wall,
Humpty Dumpty had a great fall,
All the king’s horses and all the king’s men,
Couldn’t put Humpty together again."

The king's men need to haul away Humpty's pieces to clear the road. Each of the king's horses can carry 50 rocks per load. If Humpty weighs 701 eggdrops, how many horses are needed to remove all of Humpty Dumpty? Note that there are 29 eggdrops per whistle, 284 grams per whistle, and 47 rocks per kilogram.
Answer #1

weight of humpty = 701 eggdrops

= 701/29 whistles { 29 eggdrops per whistle }

= 284*701/29 grams { 284 grams per whistle }

= 7657.077 grams

= 7.657077 kg {1 kg = 1000 grams}

=  47 * 7.657077 rocks { 47 rocks per kilogram }

= 359.88 rocks

no. of horses required = (total load) / (load 1 horse can carry)

= 359.88 / 50

= 7.1976

since no. of horses can be integer,

so, no. of horses required = 8

P.S. (please upvote if you find the answer satisfactory)